# Nuisance parameters and elliptically symmetric distributions: a geometric approach to parametric and semiparametric efficiency

### v2

@@ VI-A Some preliminaries
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}=_{d}-2^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})+\varphi_{0}(\mathcal{Q})\left(\sqrt{\mathcal{Q}}\mathbf{u}^{T}\bm{\Sigma}_{0}^{-1/2}\bm{\mu}_{i}^{0}+2^{-1}\mathcal{Q}\mathbf{u}^{T}\mathbf{P}_{i}^{0}\mathbf{u}\right),\;i=1,\ldots,d.\end{split}$ (110)
Consequently, the entries of the $q$-dimensional score vector of the parameters of interest can be simply expressed as:
$\begin{split}[\mathbf{s}_{\bm{\gamma}_{0}}]_{i}=[\mathbf{s}_{\bm{\theta}_{0}}]_{i}\;i=1,\ldots,q,\end{split}$ (111)
while the $r$-dimensional score vector of the finite-dimensional nuisance parameters is given by:
$\begin{split}[\mathbf{s}_{\bm{\xi}_{0}}]_{j}=[\mathbf{s}_{\bm{\theta}_{0}}]_{q+j}\;j=1,\ldots,r.\end{split}$ (112)
@@
To derive the FIM $\mathbf{I}_{\bm{\theta}_{0}}$, we may use the procedure in [27] that leads to the following compact expression given in [8, Sect. 1.6.5] as:
@@
$\begin{split}\mathbf{I}_{\bm{\theta}_{0}}=\beta(g_{0})&\mathbf{J}[\bm{\mu}_{0}]^{T}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\\ &+2^{-1}\alpha(g_{0})\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}+2^{-1}(1-\alpha(g_{0})^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]\end{split}$ (113)
@@
where the scalars $\beta(g_{0})$ and $\alpha(g_{0})$ are given in (64) and (63) respectively, while the Jacobian matrices, $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, are explicitly expressed as
@@
VI-B Conditions of equality between efficient SFIM and parametric efficient FIM
By collecting the previous results, we are ready to state the following proposition whose proof is given in Appendix .9:
@@ VI-B Conditions of equality between efficient SFIM and parametric efficient FIM
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ be a RES distributed vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})$ are parameterized by $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Gamma\times\Psi$, where $\bm{\gamma}_{0}\in\Gamma$ is the $q$-dimensional vector of interest and $\bm{\xi}_{0}\in\Psi$ is the $r$-dimensional nuisance vector, such that $d=q+r$. Let $g_{0}\in\overline{\mathcal{G}}$ be the infinite-dimensional nuisance parameter. Then, the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ for the model $\mathcal{P}_{\bm{\theta},g}$ in (102) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in $\mathcal{P}_{\bm{\theta},g_{0}}$ iff the following condition is satisfied:
@@
As examples of how condition (115) can be used, let us consider two quite common parameterization in signal processing application: the case where the location vector and the scatter matrix have no parameters in common and the “low-rank”, parameterization model.
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ is parameterized by the parameter of interest $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ while the scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$ is parameterized by the nuisance parameter $\bm{\xi}_{0}\in\Psi\subseteq\mathbb{R}^{r}$. Furthermore, we assume that such parameterisation satisfies the assumptions P1 and P2 given in Sect. VI-A.
@@
$\mathcal{P}_{\bm{\gamma},\bm{\xi},g}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g)=|\bm{\Sigma}(\bm{\xi})|^{-1/2}g((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,g\in\overline{\mathcal{G}}\right\},$ (116)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},g_{0})=|\bm{\Sigma}(\bm{\xi})|^{-1/2}g_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\},$ (117)
$\mathcal{P}_{\bm{\gamma}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi}_{0},g_{0})=|\bm{\Sigma}(\bm{\xi}_{0})|^{-1/2}g_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi}_{0})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma\right\}.$ (118)
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},g_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ and $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$. Then, the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},g_{0})$ for the model (116) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in (117). Moreover this latter FIM is equal to the FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ for the model (118).
@@
As a direct consequence of this specific parameterisation, we immediately have that $\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})]=\mathbf{0}$ and $\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]=\mathbf{0}$. Furthermore, since $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, it follows from (113) that the FIM $\mathbf{I}_{\bm{\theta}_{0}}$ is block-diagonal, i.e. $\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}=\mathbf{0}$. Consequently the condition (115) of Proposition VI.1 is satisfied and $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ reduces to $\mathbf{I}_{\bm{\gamma}_{0}}$. ∎

### v3

@@ VI-A Some preliminaries
$\begin{split}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}=_{d}-2^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})+\bar{\varphi}_{0}(\mathcal{Q})\left(\sqrt{\mathcal{Q}}\mathbf{u}^{T}\bm{\Sigma}_{0}^{-1/2}\bm{\mu}_{i}^{0}+2^{-1}\mathcal{Q}\mathbf{u}^{T}\mathbf{P}_{i}^{0}\mathbf{u}\right),\;i=1,\ldots,d.\end{split}$ (112)
where $\mathbf{P}_{i}^{0}\triangleq\bm{\Sigma}_{0}^{-1/2}\bm{\Sigma}_{i}^{0}\bm{\Sigma}_{0}^{-1/2}$, $\bm{\Sigma}_{i}^{0}\triangleq{\frac{\partial\bm{\Sigma}(\bm{\theta})}{\partial\theta_{i}}}|_{\bm{\theta}=\bm{\theta}_{0}}$ and $\bm{\mu}_{i}^{0}\triangleq\frac{\partial\bm{\mu}(\bm{\theta})}{\partial\theta_{i}}|_{\bm{\theta}=\bm{\theta}_{0}}.$ Consequently, $\mathbf{s}_{\bm{\gamma}_{0}}]_{i}=[\mathbf{s}_{\bm{\theta}_{0}}]_{i}$, $i=1,\ldots,q$ and $\mathbf{s}_{\bm{\xi}_{0}}]_{j}=[\mathbf{s}_{\bm{\theta}_{0}}]_{q+j}$, $j=1,\ldots,r$.
@@
To derive the FIM $\mathbf{I}_{\bm{\theta}_{0}}$, we may use the procedure in [37] that leads to the following compact expression given in [9, Sect. 1.6.5] as:
@@
$\begin{split}\mathbf{I}_{\bm{\theta}_{0}}=\beta(\bar{g}_{0})&\mathbf{J}[\bm{\mu}_{0}]^{T}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]\\ &+2^{-1}\alpha(\bar{g}_{0})\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}+2^{-1}(1-\alpha^{-1}(\bar{g}_{0}))\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})],\end{split}$ (113)
@@
where the scalars $\beta(\bar{g}_{0})$ and $\alpha(\bar{g}_{0})$ are given in (64) and (63) respectively, while the Jacobian matrices, $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, are explicitly expressed as
@@
We can now introduce the semiparametric efficient score vector $\bar{\mathbf{s}}_{\bm{\theta}_{0}}=\mathbf{s}_{\bm{\theta}_{0}}-\Pi(\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{T}^{c}_{\bar{g}_{0}})$ for $\bm{\theta}_{0}\in\Theta$ in the semiparametric model $\mathcal{P}_{\bm{\theta},\bar{g}}$ in (108). From the expression of $\mathbf{s}_{\bm{\theta}_{0}}$ in (112) and the one of the projection operator $\Pi(\cdot|\mathcal{T}^{c}_{\bar{g}_{0}})$ in (81), we have that:
$\begin{split}[\bar{\mathbf{s}}_{\bm{\theta}_{0}}]_{i}&=[\mathbf{s}_{\bm{\theta}_{0}}]_{i}-E\{[\mathbf{s}_{\bm{\theta}_{0}}]_{i}|\mathcal{Q}\}-E\left\{\mathcal{Q}[\mathbf{s}_{\bm{\theta}_{0}}]_{i}\right\}\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m)\\ &=\sqrt{\mathcal{Q}}\bar{\varphi}_{0}(\mathcal{Q})\mathbf{u}^{T}\bm{\Sigma}_{0}^{-1/2}\bm{\mu}_{i}^{0}+2^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})\left(\mathbf{u}^{T}\mathbf{P}_{i}^{0}\mathbf{u}-m^{-1}\mathrm{tr}(\mathbf{P}_{i}^{0})\right)-\mathrm{tr}(\mathbf{P}_{i}^{0})\sigma_{\mathcal{Q}}^{-2}(\mathcal{Q}-m),\end{split}$ (115)
where we have used the equality $E\{\mathcal{Q}^{2}\bar{\varphi}_{0}(\mathcal{Q})\}=m(m+2)$ given in footnote 3.
Moreover, by direct calculation, the efficient semiparametric FIM for $\bm{\theta}_{0}$ in $\mathcal{P}_{\bm{\theta},\bar{g}}$ can be expressed as:
$\begin{split}\bar{\mathbf{I}}_{\bm{\theta}_{0}}&=\beta(\bar{g}_{0})\mathbf{J}[\bm{\mu}_{0}]^{T}\bm{\Sigma}_{0}^{-1}\mathbf{J}[\bm{\mu}_{0}]+\\ &\quad+2^{-1}\alpha(\bar{g}_{0})\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\left[\bm{\Sigma}_{0}^{-1}\otimes\bm{\Sigma}_{0}^{-1}+(2\alpha^{-1}(\bar{g}_{0})\sigma_{\mathcal{Q}}^{-2}-m^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\mathrm{vec}(\bm{\Sigma}_{0}^{-1})^{T}\right]\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})],\end{split}$ (116)
where we recall that $\sigma_{\mathcal{Q}}^{2}\triangleq E\{(\mathcal{Q}-m)^{2}\}=E\{\mathcal{Q}^{2}\}-m^{2}$.  [^7] It is important to note that the expression of the efficient semiparametric score vector $\bar{\mathbf{s}}_{\bm{\theta}_{0}}$ and the related FIM $\bar{\mathbf{I}}_{\bm{\theta}_{0}}$ derived in eqs. (43) and (46) of [26] are incorrect. In fact, in the derivation proposed in [26], the projection was incorrectly taken with respect to the “unconstrained” nuisance tangent space $\mathcal{T}_{g_{0}}^{u}$, whereas $\mathbf{s}_{\bm{\theta}_{0}}$ must be projected onto the “constrained” one, i.e. $\mathcal{T}_{\bar{g}_{0}}^{c}$.
Remark: One can expect that, knowing the density generator $\bar{g}_{0}$ should lead to a larger (in the sense of the positive definite matrices ordering) FIM for $\bm{\theta}_{0}$. However, proving that ${\mathbf{I}}_{\bm{\theta}_{0}}\geq\bar{\mathbf{I}}_{\bm{\theta}_{0}}$ through the inequality $2^{-1}(1-\alpha^{-1}(\bar{g}_{0}))\geq 2\alpha^{-1}(\bar{g}_{0})\sigma_{\mathcal{Q}}^{-2}-m^{-1}$ from (113), (116) for arbirtary RES distribution seems quite involved. Nevertheless, it is straightforward to prove that for the Gaussian distribution for which $\mathcal{Q}$ is $\chi_{m}^{2}$ distributed, $\alpha(\bar{g}_{0})=1$ and $\sigma_{\mathcal{Q}}^{2}=2m$ imply ${\mathbf{I}}_{\bm{\theta}_{0}}=\bar{\mathbf{I}}_{\bm{\theta}_{0}}$ and for the $t$ distribution, $\alpha(\bar{g}_{0})=\frac{m+\nu}{m+\nu+2}$ and $\sigma_{\mathcal{Q}}^{2}=(\frac{2m}{\nu-4})(m+\nu-2)$ imply ${\mathbf{I}}_{\bm{\theta}_{0}}>\bar{\mathbf{I}}_{\bm{\theta}_{0}}$.
The calculation of the efficient semiparametric score vector of the parameters of interest $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}=\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}}+\mathcal{T}^{c}_{\bar{g}_{0}})$ and the related efficient semiparametric FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{0})$ is more difficult task. In fact, even if the projection $\Pi(h|\mathcal{T}_{\bm{\xi}_{0}}+\mathcal{T}^{c}_{\bar{g}_{0}})$, $\forall h\in\mathcal{H}$ can be explicitly obtained as shown in Appendix .4 (see also [1, Theo. 5, A.4]), the calculation may be quite involved depending on the parameterisation at hands.
It may therefore be of practical interest to ask the following question: is it possible to characterize all parameterisations $\bm{\theta}\mapsto(\bm{\mu}(\bm{\theta}),\bm{\Sigma}(\bm{\theta}))$ that imply that not knowing the finite-dimensional nuisance vector $\bm{\xi}_{0}\in\Psi$, while knowing $\bar{g}_{0}\in\overline{\mathcal{G}}$, leads to the same loss of efficiency as not knowing $\bm{\xi}_{0}\in\Psi$ nor the density generator $\bar{g}_{0}$?
As we will see below, many parameterisations used in applications satisfy this adaptivity property. We will therefore be able to easily compute the relevant $\bar{\mathbf{s}}_{\bm{\gamma}_{0}}$ and $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{0})$.
VI-B An adaptivity condition for the parameterized case
By collecting the previous results, we are ready to state the following original proposition:
@@ VI-B An adaptivity condition for the parameterized case
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},\bar{g}_{0})$ be a RES distributed vector whose location vector $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\theta}_{0})$ and scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\theta}_{0})$ are parameterized by $\bm{\theta}_{0}=(\bm{\gamma}_{0}^{T},\bm{\xi}_{0}^{T})^{T}\in\Gamma\times\Psi$, where $\bm{\gamma}_{0}\in\Gamma$ is the $q$-dimensional vector of interest and $\bm{\xi}_{0}\in\Psi$ is the $r$-dimensional nuisance vector, such that $d=q+r$. Let $\bar{g}_{0}\in\overline{\mathcal{G}}$ be the infinite-dimensional nuisance parameter. Then, the efficient SFIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{0})$ for the model $\mathcal{P}_{\bm{\theta},\bar{g}}$ in (108) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in $\mathcal{P}_{\bm{\gamma},\bm{\xi}}$ in (109) iff the following condition is satisfied:
@@
Proof:
Let us apply Lemma II.2 to the case of RES distributions, we can identify $\mathcal{T}_{2}$ with $\mathcal{T}_{\bm{\xi}_{0}}$ in (111) and $\mathcal{T}_{3}$ with $\mathcal{T}^{c}_{\bar{g}_{0}}$ in (79). Then, Condition (28) of Lemma II.2 can be expressed as:
$\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}})]\perp\mathcal{T}^{c}_{\bar{g}_{0}}.$ (118)
Since, as shown in Appendix .8, $\mathcal{T}^{u}_{g_{0}}=\mathcal{T}^{c}_{\bar{g}_{0}}\oplus\mathrm{Span}\{\mathcal{Q}-m\}$, a necessary and sufficient condition for (118) to hold is that:
$\Pi\left(\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}})|\mathcal{T}^{u}_{g_{0}}\right)]_{i}\in\mathrm{Span}\{\mathcal{Q}-m\},$ (119)
for $i=1,\ldots,q$, that is, from (78), $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}-\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}})|\mathcal{Q}\right\}]_{i}\in\mathrm{Span}\{\mathcal{Q}-m\}$
By using the expression of $\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}})$ deduced from (14), we have:
$E\left\{\Pi(\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{T}_{\bm{\xi}_{0}})|\mathcal{Q}\right\}=\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}\mathbf{I}_{\bm{\xi}_{0}}^{-1}E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\}.$ (120)
By definition, $E\left\{\mathbf{s}_{\bm{\gamma}_{0}}|\mathcal{Q}\right\}$ and $E\left\{\mathbf{s}_{\bm{\xi}_{0}}|\mathcal{Q}\right\}$ are the two sub-vector of $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$, then, to conclude the proof, we just need to evaluate $E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}$. To this end, from the expression of $\mathbf{s}_{\bm{\theta}_{0}}$ given in (112) and by noticing that:
we immediately have that:
$\begin{split}E\{\mathbf{s}_{\bm{\theta}_{0}}|\mathcal{Q}\}&=_{d}2^{-1}(m^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-1)\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1})\\ &=2^{-1}(m^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-1)[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]^{T}\mathrm{vec}(\bm{\Sigma}_{0}^{-1}).\end{split}$ (122)
Consequently, from (122) and (120), the condition (119) holds true if and only if
$a_{i}2^{-1}(m^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})-1)\in\mathrm{Span}\{\mathcal{Q}-m\},$ (123)
for $i=1,\ldots,q$ and where $a_{i}$ are the $q$ components of the vector in the left-hand side of (117).
To conclude the proof, we can trivially verified that  [^8] Note that $\mathrm{Pr}\left(\{1+m^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q})=0\}\right)=0$, since $\mathcal{Q}$ is a continuous random variable. $(1+m^{-1}\mathcal{Q}\bar{\varphi}_{0}(\mathcal{Q}))\notin\mathrm{Span}\{(\mathcal{Q}-m)\}$. Consequently, (123) is verified if and only if $a_{i}=0$, for all $i=1,\ldots,q$ that is equivalent to the condition in (117). ∎
As examples of how condition (117) can be used, let us consider two quite common parametrizations in signal processing application: the case where the location vector and the scatter matrix have no parameters in common and the “low-rank”, parameterization model.
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},\bar{g}_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ is parameterized by the parameter of interest $\bm{\gamma}_{0}\in\Gamma\subseteq\mathbb{R}^{q}$ while the scatter matrix $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$ is parameterized by the nuisance parameter $\bm{\xi}_{0}\in\Psi\subseteq\mathbb{R}^{r}$. Furthermore, we assume that such parameterisation satisfies the assumptions P1 and P2 given in Sect. VI-A.
@@
$\mathcal{P}_{\bm{\gamma},\bm{\xi},\bar{g}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},\bar{g})=|\bm{\Sigma}(\bm{\xi})|^{-1/2}\bar{g}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi,\bar{g}\in\overline{\mathcal{G}}\right\},$ (124)
$\mathcal{P}_{\bm{\gamma},\bm{\xi}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi},\bar{g}_{0})=|\bm{\Sigma}(\bm{\xi})|^{-1/2}\bar{g}_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma,\bm{\xi}\in\Psi\right\},$ (125)
$\mathcal{P}_{\bm{\gamma}}\triangleq\left\{p_{X}(\mathbf{x}|\bm{\gamma},\bm{\xi}_{0},\bar{g}_{0})=|\bm{\Sigma}(\bm{\xi}_{0})|^{-1/2}\bar{g}_{0}((\mathbf{x}-\bm{\mu}(\bm{\gamma}))^{T}\bm{\Sigma}(\bm{\xi}_{0})^{-1}(\mathbf{x}-\bm{\mu}(\bm{\gamma}))):\bm{\gamma}\in\Gamma\right\}.$ (126)
@@
Let $\mathbb{R}^{m}\ni\mathbf{x}\sim RES_{m}(\bm{\mu}_{0},\bm{\Sigma}_{0},\bar{g}_{0})$ where $\bm{\mu}_{0}\triangleq\bm{\mu}(\bm{\gamma}_{0})$ and $\bm{\Sigma}_{0}\triangleq\bm{\Sigma}(\bm{\xi}_{0})$. Then, the efficient Semiparametric FIM (SFIM) $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0},\bar{g}_{0})$ for the model (124) is equal to the parametric efficient FIM $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ in (125). Moreover this latter FIM is equal to the FIM $\mathbf{I}_{\bm{\gamma}_{0}}$ for the model (126).
@@
For this specific parameterisation, we trivially have that $\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})]=\mathbf{0}$ and $\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]=\mathbf{0}$. Furthermore, since $\mathbf{J}[\bm{\mu}_{0}]=[\mathbf{J}_{\bm{\gamma}}[\bm{\mu}_{0}],\mathbf{J}_{\bm{\xi}}[\bm{\mu}_{0}]]$ and $\mathbf{J}[\mathrm{vec}(\bm{\Sigma}_{0})]=[\mathbf{J}_{\bm{\gamma}}[\mathrm{vec}(\bm{\Sigma}_{0})],\mathbf{J}_{\bm{\xi}}[\mathrm{vec}(\bm{\Sigma}_{0})]]$, it follows from (113) that the FIM $\mathbf{I}_{\bm{\theta}_{0}}$ is block-diagonal, i.e. $\mathbf{I}_{\bm{\gamma}_{0}\bm{\xi}_{0}}=\mathbf{0}$. Consequently the condition (117) of Proposition VI.1 is satisfied and $\bar{\mathbf{I}}(\bm{\gamma}_{0}|\bm{\xi}_{0})$ reduces to $\mathbf{I}_{\bm{\gamma}_{0}}$. ∎